x+y=38
x=2y-25
must solve using substitution

Answer by checkley71(8403)

IF X=2Y-25 THEN SUBSTITUTING (2Y-25) FOR X IN EQUATION 1 WE GET
(2Y-25)+Y=38 OR 3Y=63 OR Y=63/3 OR Y=21 THE X=2*21-25 OR X=42-25 OR X=17
PROOF 1+21=38 & 17=2*21-25OR 17=42-25 OR 17=17